OpenStudy (anonymous):

If g(x) = 3(x - 2), find the value of x if g(x) = 6. 12 2.6 8 4 I Got D @texaschic101

OpenStudy (texaschic101):

g(x) = 3(x - 2) 6 = 3(x - 2) 6 = 3x - 6 6 + 6 = 3x 12 = 3x 4 = x yep..right again

OpenStudy (anonymous):

Shanks!!!

OpenStudy (texaschic101):

lol....sure thing